Find an equation for the line tangent to the graph of y=x3+ √xy + y3=3 at the point (1,1).
Answer #1

SOLUTION :


Curve : x^3 + sqrt(xy) + y^3 = 3

=> x^3 + y^3 + x^(1/2) * y^(1/2) = 3


Differentiate w.r.to x :

=> 3x^2 + 3y^2 dy/dx - (1/(2x^(1/2)  * y^(1/2) + x^(1/2) / (2y^(1/2) dy/dx) = 0

=> (3y^2 - 1/2 (x/y)^(1/2)) dy/dx = (-3x^2 + 1/2 (y/x)^(1/2))

=> dy/dx = (-3x^2 + 1/2 (y/x)^(1/2) / (3y^2 - 1/2 (x/y)^(1/2))


So, slope of tangent at point (1.1) of the curve 

= (-3 + 1/2 * 1)/ (3 - 1/2 * 1)

= (-5/2) / (5/2)

= - 1 


Equation of this tangent through point (1, 1) :

=> (y - 1) = - 1 (x - 1) (point-slope form (y- y1) = m(x - x1))

=> y - 1 = - x + 1

=> y = - x + 2  (ANSWER)

answered by: Tulsiram Garg